# A Method for Auto-Differentiation of the Voronoi Tessellation

### v2

@@ Abstract
Voronoi tessellation, also known as Voronoi diagram, is an important computational geometry technique that has applications in various scientific disciplines. It involves dividing a given space into regions based on the proximity to a set of points. Autodifferentiation is a powerful tool for solving optimization tasks. Autodifferentiation assumes constructing a computational graph that allows to compute gradients using backpropagation algorithm. However, often the Voronoi tessellation remains the only non-differentiable part of a pipeline, prohibiting end-to-end differentiation. We present the method for autodifferentiation of the 2D Voronoi tessellation. The method allows one to construct the Voronoi tessellation and pass gradients, making the construction end-to-end differentiable. We provide the implementation details and present several important applications. To the best of our knowledge this is the first autodifferentiable realization of the Voronoi tessellation providing full set of Voronoi geometrical parameters in a differentiable way.
Machine Learning, ICML
@@ 1 Introduction
The versatility of Voronoi diagrams extends to fields such as physics (Kobayashi & Sugihara, 2002; de Lacy Costello et al., 2004; Debnath et al., 2016), climate modeling (Ju et al., 2011; Liu et al., 2023), biology (Jungck et al., 2021; Marcelpoil & Usson, 1992; Edla & Jana, 2012), material science (Sun et al., 2018; Acton et al., 2018), and geographic information systems (GIS) (Nowak, 2015), providing crucial insights in each domain. For instance, in biology, they help model cellular growth patterns, while in GIS, they aid in efficient land use planning. Despite their widespread utility, the application of Voronoi tessellation in optimization, especially in solving inverse problems such as those encountered in fluid dynamics, presents unique challenges. These problems often require the integration of observational data, to optimize the placement of reference points for numerical simulations. Such problems can be solved efficiently using optimization algorithms and machine learning. There are a large number of optimization algorithms, some of the most efficient of which are differentiable optimization methods. However, obtaining Voronoi tessellation characteristics, such as Voronoi region areas or adjacent edge lengths, involves the use of Voronoi tessellation algorithms, which are non-differentiable. In this paper, we introduce a novel method that enables automatic differentiation of the process of construction of the Voronoi tessellation. This approach promises to enhance the solution of inverse problems across various disciplines by bridging the gap between the differentiable optimization methods and the non-differentiable nature of the Voronoi tessellation algorithm.
@@ 1.1 Related Work
There are several works that are considered related. In (Chen et al., 2022) the authors combined Voronoi tessellation with normalizing flows to construct a new type of invertible transformation. This transformation facilitates applications such as Voronoi dequantization and disjoint mixture models, where discrete values are mapped into a continuous space. However, the method described therein does not make broad range of the Voronoi tessellation parameters differentiable w.r.t. input of Voronoi tessellation (point clouds). Instead, the authors focused on integrating Voronoi tessellation into a framework for modeling the transition between discrete and continuous spaces. This integration uses Voronoi tessellation to partition space and create mappings that are amenable to gradient-based optimization. However, the method described in the paper does not inherently support the differentiation of the parameters of Voronoi tessellations, such as region areas, with respect to the input of the tessellation (e.g., point cloud). The tessellation is used as a tool within a larger framework, but the process of tessellation itself, including the derivation of specific parameters like cell volumes, is not set up for direct differentiation in relation to the input data.
@@ 2.1 Voronoi tessellation
The input of our method is the set of points $S$ with more than three points. We formally define Voronoi region for a given point as $VR(p,S)$ or $VR(p)$ if it’s clear what $S$ is mentioned. $VR(p)$ consists of all points in $\mathbf{R}^{2}$ for which $p$ is the nearest neighbour site.
@@ 2.2 Delaunay triangulation
The Delaunay triangulation has a wide range of applications in various fields, from computer graphics and image processing to geographic information systems and computational biology. It is commonly used for surface reconstruction, mesh generation, terrain modeling, and Voronoi diagram computation. It is also used in finite element analysis, where it provides a more accurate and efficient representation of the geometry. In addition, Delaunay triangulation is used in pattern recognition, data analysis, and machine learning algorithms, where it helps to identify clusters and patterns in large datasets. We define Delaunay triangulation of a set of site points $S$ as $DT(S)$.
@@ 3 Auto-differentiable Voronoi tessellation
First we obtain the $DT(S)$. We do not include the calculation of $DT(S)$ in the computational graph of AD. We construct the computational graph dynamically based on the adjacency information obtained after $DT(S)$. Therefore, to compute Delaunay triangulation we may choose any available software. Because of duality between $DT(S)$ and $V(S)$ we may calculate the geometry of $V(S)$ based on the pre-computed $DT(S)$.

### v3

@@ Abstract
We present the method for autodifferentiation of the 2D Voronoi tessellation. The method allows one to construct the Voronoi tessellation and pass the gradients, making any numerical construction involving it end-to-end differentiable. We provide the implementation details and test it on several problems, such as optimal hospital placement. To the best of our knowledge this is the first autodifferentiable realization of the Voronoi tessellation providing full set of Voronoi geometrical parameters in a differentiable way.
Machine Learning
@@ 1 Introduction
The versatility of Voronoi diagrams extends to fields such as physics (Kobayashi & Sugihara, 2002; de Lacy Costello et al., 2004; Debnath et al., 2016), climate modeling (Ju et al., 2011; Liu et al., 2023), biology (Jungck et al., 2021; Marcelpoil & Usson, 1992; Edla & Jana, 2012), material science (Sun et al., 2018; Acton et al., 2018), and geographic information systems (GIS) (Nowak, 2015), providing crucial insights in each domain. For instance, in biology, they help model cellular growth patterns, while in GIS, they aid in efficient land use planning. Despite their widespread utility, the application of Voronoi tessellation in optimization, especially in solving inverse problems such as those encountered in fluid dynamics, presents unique challenges. These problems often require the integration of observational data, to optimize the placement of reference points for numerical simulations. Such problems can be solved efficiently using optimization algorithms and machine learning. There are a large number of optimization algorithms, some of the most efficient of which are differentiable optimization methods. However, obtaining Voronoi tessellation characteristics, such as Voronoi region areas or adjacent edge lengths, involves the use of Voronoi tessellation algorithms, which are non-differentiable.
In this paper, we introduce a novel method that enables automatic differentiation of the process of construction of the Voronoi tessellation. This approach promises to enhance the solution of inverse problems across various disciplines by bridging the gap between the differentiable optimization methods and the non-differentiable nature of the Voronoi tessellation algorithm.
@@ 1.1 Related Work
There are several works that are considered related. The work (diff_delaunay) presents a method to get a differentiable version of the Delaunay triangulation. In (Chen et al., 2022) the authors combined Voronoi tessellation with normalizing flows to construct a new type of invertible transformation. This transformation facilitates applications such as Voronoi dequantization and disjoint mixture models, where discrete values are mapped into a continuous space. However, the method described therein does not make broad range of the Voronoi tessellation parameters differentiable w.r.t. input of Voronoi tessellation (point clouds). Instead, the authors focused on integrating Voronoi tessellation into a framework for modeling the transition between discrete and continuous spaces. This integration uses Voronoi tessellation to partition space and create mappings that are amenable to gradient-based optimization. However, the process of tessellation itself, including the derivation of specific parameters like cell volumes, is not set up for direct differentiation in relation to the input data.
@@ 2.1 Voronoi tessellation
The input of our method is the set of points $S$ with more than three points. We formally define Voronoi region for a given point $p$ as $VR(p,S)$ or $VR(p)$ if it’s clear what $S$ is mentioned. $VR(p)$ consists of all points in $\mathbf{R}^{2}$ for which $p$ is the nearest neighbour site.
@@ 2.2 Delaunay triangulation
Delaunay triangulation is a pivotal element of our method, as it plays a crucial role in the computation of Voronoi diagrams.
We define Delaunay triangulation of a set of site points $S$ as $DT(S)$.
@@ 3 Auto-differentiable Voronoi tessellation
For simplicity, we assume that $S$ are in general position. First, we obtain the $DT(S)$. We do not include the calculation of $DT(S)$ in the computational graph of AD. We construct the computational graph dynamically based on the adjacency information obtained after $DT(S)$. Therefore, to compute Delaunay triangulation we may choose any available software. Because of the duality between $DT(S)$ and $V(S)$ we may calculate the geometry of $V(S)$ based on the pre-computed $DT(S)$.